### v1

@@ 3.   SAR with error spatial dependence
where the $v_{j}$ are $i.i.d$ with mean 0 and variance 1, and the $b_{ij}\equiv b_{ijn}$ can depend on $n$. This is the Kelejian and Prucha (2007) approach to error spatial dependence, and indeed we will employ the spatial heteroskedasticity and autocorrelation consistent (SHAC) method pioneered in that paper. Our imposition that $z_{i}$ be uncorrelated with $\epsilon_{i}$ now, and subsequently, requires that $z_{i}$ is uncorrelated with all the $v_{j}$, $j=1,\ldots,n$.
@@ 3.1   Test statistic
$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel function that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
Assumption 7 restricts the spatial dependence in the errors to a manageable degree, see e.g. Kelejian and Prucha (2007), Delgado and Robinson (2015) and Gupta and Qu (2024) for similar assumptions. Assumption 8 is a standard assumption on kernels in the HAC setting, see e.g. Kelejian and Prucha (2007).
@@
$\hat{\Xi}_{1,rs}=n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{ri}k_{sj}\hat{u}_{j}\hat{u}_{j}\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),r,s=1,\ldots,J_{1}.$ (3.3)
@@ 3.2   Asymptotic properties
In all subsequent lemmas and theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9. In Lemma 3.1 and Theorems 3.1-3.2, which follow, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$. Furthermore, for any matrix $A$, let $\|A\|=\left\{\overline{\alpha}\left(A^{\prime}A\right)\right\}^{\frac{1}{2}}$ i.e. the spectral norm of $A$.
@@
Then,
$\big{\|}\hat{\Xi}_{1}-\Xi_{1}\big{\|}=o_{p}(1).$
The condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $d_{\xi}/n\rightarrow 0$ by Assumption 9(a). Now, define $\mathcal{M}_{2}=\frac{1}{n}B^{\prime}K_{1}\mathcal{V}_{2}K_{1}^{\prime}B$, with $\mathcal{V}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$. Then we have the following theorem.
@@
To establish the asymptotic distribution of $\mathbb{W}_{2}$, the following assumption is imposed to check a Lyapunov condition, as also e.g. by Delgado and Robinson (2015) and Gupta and Qu (2024).
@@
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{2}$ provides a consistent test.
[…]
Observe that the condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies $\frac{d_{\xi}^{3}}{n}\rightarrow 0$ if $\eta\geq 1/3$, so depending on the value of $q$, the rate condition (3.5) can be simplified.
@@ 4.   Misspecification robustness
We now consider a degree of misspecification robustness in (2.1), which allows $\epsilon_{i}$ to have the structure in Section 3. Focusing on the case $d_{\lambda}=1$ for notational simplicity, and writing $W_{1}=W$, (2.1) imposed the parametric form $\lambda Wy$ and took $W$ as known in the “spatial lag” term. We now allow the spatial lag in outcomes to take the nonparametric form $Gy$, where $G$ has elements $g\left(d_{rs}\right)$, $r,s=1,\ldots,n$, for some vector of exogenous (independent of $v_{j}$, $j=1,\ldots,n$) economic distance measures $d_{rs}$. This idea was introduced by Pinkse et al. (2002), and has been much used since (see e.g. Sun, 2016 and Gupta and Qu, 2024).
@@ 4.1   Test statistic
We approximate the elements $g\left(d_{ij}\right)$ of $G$ with a series of basis functions. Following Pinkse et al. (2002), we suppose
@@
where $\tau_{l}$ are unknown coefficients, and the $e_{l}$ form a basis of the function space to which $g$ belongs. Let $c_{i}^{\prime}$ be the $i$-th row of the $n\times d_{\tau}$ matrix $\mathfrak{C}$ with typical $(i,l)$-th element $\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $\tau=\left(\tau_{1},\ldots,\tau_{d_{\tau}}\right)^{\prime}$, $d_{\tau}\rightarrow\infty$ as $n\rightarrow\infty$. Then we generalize (2.1) to
@@
and (2.3) to
@@
with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$. In the matrix notation, we now have
@@
where $F=[\mathfrak{C},X,\Psi]$ and $\theta=\left(\tau^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. Our test is implemented using the 2SLS estimator
@@ 4.2   Asymptotic properties
Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix. The following assumption generalizes Assumption 5 to the setting of this section.
@@
$\left(I_{n}-G\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely, uniformly in $d_{rs},r,s=1,\ldots,n$.
@@
as $n\rightarrow\infty$, the following hold:
[…]
(2) $\mathbb{W}_{3}$ provides a consistent test.
[…]
@@ 5.   Varying spatial coefficients
In this section we revert to the case of known $W_{j},j=1,\ldots,d_{\lambda}$, but now allow the spatial lag parameters $\lambda_{j}$ to vary. The model we consider is closely related to Malikov and Sun (2017) and Sun and Malikov (2018), with the difference that we do not study panel data settings. Instead, we allow $d_{\beta}\rightarrow\infty$ and general spatial dependence in the disturbances. Our model is also related to the one studied by Sun (2024), wherein the model includes a nonparametric nuisance regression function but does not feature dependent disturbances. Thus, our model complements the extant literature.
@@ 5.1   Test statistic
with $\lambda_{j}(\cdot)$ unknown scalar-valued functions and $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j}$, i.e. error with spatial dependence. Introduce the approximation $\lambda_{j}\left(z_{i}\right)=\sum_{k=1}^{l_{j}}\mu_{kj}\phi_{kj}\left(z_{i}\right)+r_{ij\lambda},j=1,\ldots,d_{\lambda}$. Set $d_{\mu}=\sum_{j=1}^{d_{\lambda}}l_{j}$, $\phi_{j}=\left(\phi_{1j},\ldots,\phi_{l_{j}j}\right)^{\prime}$, and let $H(z)$ be the $n\times d_{\mu}$ matrix with $i$-th row $\left(y^{\prime}w_{i,1}\phi_{1}\left(z_{i}\right),\ldots,y^{\prime}w_{i,d_{\lambda}}\phi_{d_{\lambda}}\left(z_{i}\right)\right)$.
Then,
@@
where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$. Defining $\Lambda_{j}(z)=diag\left(\lambda_{j}\left(z_{1}\right),\ldots,\lambda_{j}\left(z_{n}\right)\right)$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
@@
where $\vartheta$ is either $\mu$ or $\alpha$ and our test will be implemented using the 2SLS estimator
@@
where $P_{K_{3}}=K_{3}\left(K_{3}^{\prime}K_{3}\right)^{-1}K_{3}^{\prime}$ with $K_{3}$ being an $n\times J_{3}$ instrument matrix with $J_{3}\geq d_{\gamma}=d_{\mu}+d_{\beta}+d_{\alpha}$, but with $J_{3}$ and $d_{\gamma}$ having the same asymptotic order. Note that we have omitted the $z$ argument for brevity.
Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. Then, the test statistic is
@@
where $d_{\vartheta}$ is either $d_{\mu}$ or $d_{\alpha}$.
Denote the elements of $G$ as $g_{ri}(z)$. We have the following assumptions:
@@
$n^{-1}K_{3}^{\prime}K_{3}$ and $\sup_{z}n^{-1}G(z)^{\prime}K_{3}K_{3}^{\prime}G(z)$ have 1.
@@
Then, we have
$\bigg{\|}\hat{\Xi}_{3}-\Xi_{3}\bigg{\|}=o_{p}(1).$
@@ 5.2   Asymptotic properties
as $n\rightarrow\infty$, the following hold:
[…]
(ii) $\mathbb{W}_{4}$ provides a consistent test.
[…]

### v2

@@ 3.  SAR with error spatial dependence
where the $v_{j}$ are $i.i.d$ with mean 0 and variance 1, and $b_{ij}\equiv b_{ijn}$ can depend on $n$. This is the Kelejian and Prucha (2007) approach to error spatial dependence, and indeed we will employ the spatial heteroskedasticity and autocorrelation consistent (SHAC) method pioneered in that paper. This type of linear process assumption is widely employed in the literature, see e.g. Robinson (2011), Robinson and Thawornkaiwong (2012), Hidalgo and Schafgans (2017) and Conley et al. (2023).
@@ 3.1 Test statistic
$\mathscr{K}(\cdot):\mathbb{R}\rightarrow[-1,1]$ is a kernel that satisfies $\mathscr{K}(0)=1$, $\mathscr{K}(x)=\mathscr{K}(-x)$, $\mathscr{K}(x)=0$ for $|x|>1$ and $|\mathscr{K}(x)-1|\leq C|x|^{\varrho},|x|\leq 1$, for some $\varrho\geq 1$.
Assumption 7 restricts the spatial dependence in the errors to a manageable degree, see e.g. Kelejian and Prucha (2007) and Delgado and Robinson (2015) for similar assumptions. Assumption 8 is a standard assumption on kernels in the HAC setting, see e.g. Kelejian and Prucha (2007).
@@
$\hat{\Xi}_{1,rs}=n^{-1}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{ri}k_{sj}\hat{u}_{i}\hat{u}_{j}\mathscr{K}\left(\min_{m}\left\{d^{*}_{ij,m}/d_{m}\right\}\right),r,s=1,\ldots,J_{1},$ (3.3)
where $\hat{u}_{i}$ are elements of the estimation residual vector $\hat{u}=y-L\hat{\xi}$.
@@ 3.2 Asymptotic properties
In all subsequent lemmas/theorems, $q$ is defined in Assumption 2 and $\eta$ is defined in Assumption 9. In Lemma 3.1 and Theorems 3.1-3.2, $d_{\xi}$ is defined as $d_{\xi}=d_{\lambda}+d_{\beta}+d_{\alpha}$. Furthermore, for any matrix $A$, let $\|A\|=\left\{\overline{\alpha}\left(A^{\prime}A\right)\right\}^{\frac{1}{2}}$ i.e. the spectral norm of $A$.
@@
Then, $\big\|\hat{\Xi}_{1}-\Xi_{1}\big\|=o_{p}(1)$.
Define $\mathcal{M}_{2}=\frac{1}{n}B^{\prime}K_{1}\mathcal{V}_{2}K_{1}^{\prime}B$, with $\mathcal{V}_{2}=(K_{1}^{\prime}K_{1})^{-1}K_{1}^{\prime}LT_{1}^{-1}R^{\prime}\mathscr{D}_{2}^{-1}RT_{1}^{-1}L^{\prime}K_{1}(K_{1}^{\prime}K_{1})^{-1}$. We have the following theorem.
@@
To derive the asymptotic distribution of $\mathbb{W}_{2}$, we need the following assumption, as also in Delgado and Robinson (2015).
@@
as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{2}\overset{d}{\rightarrow}N(0,1)$. (ii) $\mathbb{W}_{2}$ provides a consistent test. (iii) Under the sequence of local alternatives $H_{\ell 2}$, $\mathbb{W}_{2}\overset{d}{\rightarrow}N(2^{-1/2},1)$.
Observe that the condition $n^{\eta-1}d_{\xi}^{2}\rightarrow 0$ implies ${d_{\xi}^{3}}/{n}\rightarrow 0$ if $\eta\geq 1/3$, so depending on the value of $q$, the rate condition (3.5) can be simplified.
@@ 4.  Misspecification robustness
Consider a degree of misspecification robustness in (2.1), where $\epsilon_{i}$ has the structure in Section 3. Focusing on the case $d_{\lambda}=1$ for notational simplicity, and writing $W_{1}=W$, (2.1) imposed the parametric form $\lambda Wy$ and took $W$ as known in this ``spatial lag'' term. We now allow this to take the nonparametric form $Gy$, where $G$ has elements $g\left(d_{rs}\right)$, $r,s=1,\ldots,n$, for some unknown function $g(\cdot)$, and a vector of exogenous (independent of $v_{j}$, $j=1,\ldots,n$) economic distance measures $d_{rs}$. This idea was introduced by Pinkse et al. (2002), and has been much used since (see e.g. Sun, 2016 and Gupta and Qu, 2024).
@@ 4.1 Test statistic
Following Pinkse et al. (2002), we approximate $g\left(d_{ij}\right)$ with a series of basis functions
@@
where $\tau_{l}$ are unknown coefficients, and $e_{l}$ form a basis of the function space to which $g(\cdot)$ belongs. Let $c_{i}^{\prime}$ be the $i$-th row of the $n\times d_{\tau}$ matrix $\mathfrak{C}$ with typical $(i,l)$-th element $\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $\tau=\left(\tau_{1},\ldots,\tau_{d_{\tau}}\right)^{\prime}$, $d_{\tau}\rightarrow\infty$ as $n\rightarrow\infty$. Then (2.1) is extended to
@@
where $g_{in}\equiv g_{i}$ is the $i$-th row of $G$, and, once again dropping $n$-subscripting, accordingly extend (2.3) to
@@
with $u_{i}=r_{iG}+r_{i\delta}+\epsilon_{i}$, where $r_{iG}=\sum_{l=d_{\tau}+1}^{\infty}\tau_{l}\sum_{j\neq i}e_{l}\left(d_{ij}\right)y_{j}$ and $r_{i\delta}=p_{i}^{\prime}\delta\left(z_{i}\right)-\psi_{i}^{\prime}\alpha$. In matrix notation, we now have
@@
where $F=[\mathfrak{C},X,\Psi]$ and $\theta=\left(\tau^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$. Our test is based on the 2SLS estimator
@@ 4.2 Asymptotic properties
Define $H_{\ell 3}\equiv H_{\ell 3,n}:\alpha=\alpha^{*}\equiv\nu_{3n}d_{\alpha}^{\frac{1}{4}}/(n\nu_{3n}^{\prime}\Gamma_{3n}\nu_{3n})^{\frac{1}{2}}$, with $\nu_{3n}$ a $d_{\alpha}\times 1$ non-zero vector and $\Gamma_{3n}$ a $d_{\alpha}\times d_{\alpha}$ matrix defined in the proofs. The following assumption generalizes Assumption 5.
@@
$\left(I_{n}-G\right)^{-1}$ exists and is uniformly bounded in row and column sums for all sufficiently large $n$, almost surely and uniformly over the support of the distance measure vector.
@@
as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N(0,1)$. (ii) $\mathbb{W}_{3}$ provides a consistent test. (iii) Under the sequence of local alternatives $H_{\ell 3}$, $\mathbb{W}_{3}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.
@@ 5.  Varying spatial coefficients
Now we revert to the case of known $W_{j},j=1,\ldots,d_{\lambda}$, but allow spatial lag parameters $\lambda_{j}$ to vary. The model is closely related to Malikov and Sun (2017) and Sun and Malikov (2018), with the difference that we do not study panel data settings. Instead, we allow $d_{\beta}\rightarrow\infty$ and general spatial dependence in the errors. Our model also relates to Sun (2024), wherein the model includes a nonparametric nuisance regression function but does not feature dependent errors. Thus, our model complements the extant literature.
@@ 5.1 Test statistic
with $\lambda_{j}(\cdot)$ unknown scalar-valued functions and $\epsilon_{i}=\sum_{j=1}^{n}b_{ij}v_{j}$, i.e. error with spatial dependence. Introduce the approximation $\lambda_{j}\left(z_{i}\right)=\sum_{k=1}^{l_{j}}\mu_{kj}\phi_{kj}\left(z_{i}\right)+r_{ij\lambda},j=1,\ldots,d_{\lambda}$. Set $d_{\mu}=\sum_{j=1}^{d_{\lambda}}l_{j}$, $\phi_{j}=\left(\phi_{1j},\ldots,\phi_{l_{j}j}\right)^{\prime}$, and let $H(z)$ be the $n\times d_{\mu}$ matrix with $i$-th row $\left(y^{\prime}w_{i,1}\phi_{1}\left(z_{i}\right),\ldots,y^{\prime}w_{i,d_{\lambda}}\phi_{d_{\lambda}}\left(z_{i}\right)\right)$. Then,
@@
where $G(z)=\left[H(z),X,\Psi\right]$, $\gamma=\left(\mu^{\prime},\beta^{\prime},\alpha^{\prime}\right)^{\prime}$, $\mu=\left(\mu_{1}^{\prime},\ldots,\mu_{d_{\mu}}^{\prime}\right)$, and $u$ has $i$-th element $u_{i}=r_{i\lambda}+r_{i\delta}+\epsilon_{i}$ with $r_{i\lambda}=\sum_{j=1}^{d_{\lambda}}r_{ij\lambda}w_{i,j}^{\prime}y$. Defining $\Lambda_{j}(z)=diag(\lambda_{j}\left(z_{1}\right),\ldots,\\ \lambda_{j}\left(z_{n}\right))$ and $S(z)=I-\sum_{j=1}^{d_{\lambda}}\Lambda_{j}(z)W_{j}$, under the usual invertibility conditions,
@@
where $\vartheta$ is either $\mu$ or $\alpha$ and our test will again be implemented using the 2SLS estimator
@@
where $P_{K_{3}}=K_{3}\left(K_{3}^{\prime}K_{3}\right)^{-1}K_{3}^{\prime}$ with $K_{3}$ being an $n\times J_{3}$ instrument matrix with $J_{3}\geq d_{\gamma}=d_{\mu}+d_{\beta}+d_{\alpha}$, but with $J_{3}$ and $d_{\gamma}$ having the same asymptotic order. Note that we have omitted the $z$ argument for brevity. Let $\mathscr{D}_{4}=\frac{1}{n}R\mathcal{U}_{3}R^{\prime}$ with $\mathcal{U}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\Xi_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$ and $T_{3}=\frac{1}{n}G^{\prime}{P_{K_{3}}}G$, and let $\widehat{\mathscr{D}}_{4}=\frac{1}{n}R\widehat{\mathcal{U}}_{3}R^{\prime}$ with $\widehat{\mathcal{U}}_{3}=T_{3}^{-1}G^{\prime}K_{3}(K_{3}^{\prime}K_{3})^{-1}\hat{\Xi}_{3}(K_{3}^{\prime}K_{3})^{-1}K_{3}^{\prime}LT_{3}^{-1}$, where $R=\left[I_{d_{\mu}},0_{d_{\mu}\times\left(d_{\beta}+d_{\alpha}\right)}\right]$ or $R=\left[0_{d_{\alpha}\times\left(d_{\mu}+d_{\beta}\right)},I_{d_{\alpha}}\right]$, and $\hat{\Xi}_{3}$ is the Kelejian and Prucha (2007) estimate of $\Xi_{3}=\frac{1}{n}K_{3}^{\prime}\Sigma K_{3}$, with $\Sigma=E(\epsilon\epsilon^{\prime})$. The test statistic is
@@
where $d_{\vartheta}$ is either $d_{\mu}$ or $d_{\alpha}$. Denote the elements of $G$ as $g_{ri}(z)$. We have the following assumptions:
@@
$n^{-1}K_{3}^{\prime}K_{3}$ and $n^{-1}G(z)^{\prime}K_{3}K_{3}^{\prime}G(z)$ have 1, the latter uniformly in $z\in\mathcal{Z}$.
@@
Then, we have $\big\|\hat{\Xi}_{3}-\Xi_{3}\big\|=o_{p}(1)$.
@@ 5.2 Asymptotic properties
as $n\rightarrow\infty$, the following hold: (i) Under $H_{0}$, $\mathbb{W}_{4}\overset{d}{\rightarrow}N(0,1)$. (ii) $\mathbb{W}_{4}$ provides a consistent test. (iii) Under the sequence of local alternatives $H_{\ell 4}$, $\mathbb{W}_{4}\overset{d}{\rightarrow}N\left(2^{-1/2},1\right)$.